OpenStudy (nikkimc):

is this correct? x^2 * x^3=x^5

OpenStudy (campbell_st):

thats correct multiply the same base... add the powers

OpenStudy (anonymous):

yep :)

OpenStudy (nikkimc):

thanks

OpenStudy (nikkimc):

\[{\left[ 4\div5y^3 \right]}^2=16/25y^6 is this correct\]

OpenStudy (anonymous):

yeah

OpenStudy (nikkimc):

thanks again

OpenStudy (anonymous):

\[=\frac{16}{25y^6}\]

OpenStudy (anonymous):

np